OpenStudy (anonymous):

how do i find the sum of a series from n=1 to infinity of ((3^n)+5)/ (4^n) ?

OpenStudy (safari321):

you will \[((3^{1}+5) \div(4^{1})\] \[((3+5)\div4\] \[8\div4\] \[=2\]

OpenStudy (anonymous):

i got that far but i need help finding r

OpenStudy (anonymous):

seperate to 2 pieces 3^n/4^n + 5/4^n

OpenStudy (anonymous):

right. i also separated them but i don't know how to find r after that

OpenStudy (anonymous):

first part has a = 3/4 and r = 3/4 second part has a = 5/4 and r = 1/4

OpenStudy (mertsj):

However (3^2+5)/4^2=14/16=7/8

OpenStudy (mertsj):

The sum is 4 2/3

OpenStudy (anonymous):

okay so mathwiz707, what do i do to create the sum with those two separate parts?

OpenStudy (anonymous):

first part sum = 3/4/(1 -3/4) and second has sum 5/4/(1-1/4) ... now just add the 2 parts

OpenStudy (anonymous):

oh okay. so is that 42/3 that you wrote as the answer?

OpenStudy (anonymous):

3 + 15 = 18 ...

OpenStudy (anonymous):

wait, what does ^ that have to do with this problem?

OpenStudy (anonymous):

means exponent

OpenStudy (anonymous):

when i added the two sums i got 14/3 because you add 3+ (5/3) not 3+15

OpenStudy (anonymous):

thanks for the help! appreciate it